Please use the Transfers Page for information on data transfer.
See the Modules Page
Please enter any significant events related to the AuScope VLBI project, including observations, configuration changes, hardware upgrades etc.
The humidity sensor on the MET4 isn't working and is going off for repair.
On pcfsyg as root:
pcfsyg:/etc/init.d# update-rc.d -f metclient.sh remove pcfsyg:/etc/init.d# update-rc.d -f metserver.sh remove pcfsyg:/etc/init.d# ./metserver.sh stop pcfsyg:/etc/init.d# ./metclient.sh stop
Also commented out the call to wx in /usr2/proc/station.prc
After installing the maser post refurb, and SSR installation for aircon compressor, the aircon was turned off at 0452 on Fri 3/6/16 and a large fan was turned on. 0430UT on 6/6/16 air con was turned on and large fan off and maser room closed.
The above has been done to see if the maser aircon at Ke is affecting the clock jump problem. Cheers, Brett
This is to let you know that the VLBI antenna on the 13/14th December had the following work done 1. Az bullgear old grease removed - new grease applied. 2. El jack screw cover was removed and more new grease applied to the jack screw arm. The jack screw cover itself was cleaned but is in need of replacing very soon. It would be good to get this done when Brett is over next. Also I need to check the gear box oil and or maybe change the oil with him? 3. The azimuth grease nipples inside the pedestal and the grease nipples on the antenna itself have had grease pumped in. 4. A general clean up was done inside the pedestal and the air conditioning filters removed and cleaned.
Three major repairs were finished on 2/12/15
* DBBC * IF wiring * Anemometer
DBBC: The Katherine DBBC is fully repaired as of 2/12/15. It originally failed 24/11/15, during r1715 with a total comms failure. This was eventually traced to the PCI7200 board, where the driver chip had burnt out. Replacing this board allowed Katherine to resume observations but with some apparent communication problems with the core2 units - bbc attenuations were being mis-set dur to a stuck bit in the communications line. The problem developed further after an initial inspection and no communications with either the CoMos or Core power levels could be made. The problem was eventually resolved on 1/12/15 when the PCI9111 board was replaced, but it may be that it was a connection issue instead - the PC frame is warped at Ke and it's possible to have some PCI lnes now connected through the backplane to the CPU.
IF wiring: The power level of SRCP going into CoMo D was unusually low in testing after the repair. It appears that the hardline or SMA connector was not making a good connection between the IF box and DBBC. Removing a replacing the hardline between the IF box output and DBBC radically improved the power level (~30 dB).
Anemometer: The anemometer at Katherine has suffered a fault and is reporting 1000 km/h winds. It has been turned off (using ispke) to avoid triggering spurious windstows. Until it is repaired, observers should use the weather radar at Katherine to look out for storm activity at Katherine.